Circular No. 7080

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA MUSCAE 1998 F. Farrell, Christies Beach, S. Australia, forwards the following precise

position for this object (cf. IAUC 7079), obtained on Jan. 1 at Stockport Observatory: α = 11^{h}56^{m}08^{s}.96, δ = −65^{o}34′20′′.7 (equinox 2000.0); his

image is available at http://www.assa.org.au/vstars/nvmus98.htm. P.

Holvorcem, Valinhos, Brazil, reports the following position end figures and magnitude from three CCD frames taken on Jan. 1.12 (0.20-m f/3.3 reflector): 08^{s}.90, 20′′.7; V = 9.4. G. Garradd, Loomberah, N.S.W., reports the following position end figures from a CCD frame taken on Jan. 3.730 (0.45-m f/5.4 reflector): 08^{s}.89, 21′′.2 (unfiltered mag 9.5).

Visual magnitude estimates: 1998 Dec. 31.41 UT, 10.1 (A. Jones, Nelson, New Zealand); 31.62, 10.0 (Jones); 31.671, 10.7 (Farrell); 1999 Jan. 1.562, 10.6 (Farrell); 1.688, 10.3 (Farrell); 1.781, 10.7 (A. Pearce, Nedlands, W. Australia); 2.669, 10.9 (Pearce); 3.804, 10.9 (Pearce).

SUPERNOVA 1998ey IN NGC 7080 J. McGaha reports the following precise position for this object (cf.

IAUC 7065), obtained with a 0.62-m reflector on 1998 Dec. 14.095 UT at Grasslands Observatory (near Sonoita, AZ) from 49 USNO A1.0 stars: α = 21^{h}30^{m}00^{s}.67, δ = +26^{o}43′21′′.4 (equinox 2000.0). A measurement with 19 GSC stars yields position end figures 00^{s}.72, 21′′.5 (uncalibrated, unfiltered magnitude ∼ 16.8). E. Prosperi, Larciano, Italy, provides the following position end figures and magnitude (0.25-m f/6.3 reflector) from a CCD frame taken on Dec. 26.730: 00^{s}.52, 24′′.5, R = 16.8.

COMET C/1998 M5 (LINEAR)

Ephemeris extension to IAUC 7020 (elements MPC 32866):

1998 /99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 23 18^{h}52.^{m}10 +39^{o}12.^{′}0 1.998 1.793 63.^{o}6 29.^{o}4 10.0
Jan. 2 18 57.05 +41 33.9 1.950 1.767 64.6 30.2  9.9
12 19 02.92 +44 44.8 1.885 1.750 66.8 31.1  9.8
22 19 09.57 +48 52.9 1.809 1.743 70.2 32.1  9.7
Feb. 1 19 16.83 +54 07.3 1.727 1.745 74.5 33.0  9.6
11 19 24.70 +60 36.4 1.647 1.757 79.4 33.5  9.5
21 19 33.22 +68 25.5 1.580 1.778 84.3 33.6  9.5
Mar. 3 19 42.8 +77 29.6 1.537 1.808 88.6 33.2  9.5
1999 January 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green